For his summer assignments, Tajee had 4 pages of the math with 13 problems on each page. He also had a reading quiz based on his book that was 3 pages long with12 questions on each. How many problems did Tajee solve in all?
1
Expert's answer
2019-09-24T11:37:38-0400

Total problems: 4*13 = 52.

Total questions from quiz: 3*12 = 36.

Total together: 52+36 = 88.

P. S. From the question we just know the number of tasks, but we don't know how many he solved.